OpenStudy (mtwizzle420):

Graph the function d= sqr root sign 1.5h which gives the approximate distance d in miles to the horizon where h is the height of the viewers eyes above the ground in feet. Approximatly how high above the ground must a person be if the horizon is 17.2 miles away?

OpenStudy (mtwizzle420):

@KenichiIvy

OpenStudy (anonymous):

17.2 = sqrt(1.5*h) 17.2^2 = 1.5h h = (17.2^2)/1.5

OpenStudy (anonymous):

To solve a balanced equation, we must get the variable by itself. To do that, we'll do the opposite operation. In this case, we see a sqrt1.5h To get rid of the square root, we'll do the opposite operation, squaring. Of course, to keep the equation balanced, we will have to do the same thing to both sides.

OpenStudy (mtwizzle420):

the choices are A.) 3 ft B.) 197 ft C.) 3 ft D.) 41 ft

OpenStudy (mtwizzle420):

Im not sure how to attach the graphs :/

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Choose a best response and close this post :)